Car batteries are often rated in amp-hours. This rating indicates the quantity of charge that can pass through the battery before the battery's chemical energy must be restored by the car's generator or some other source. The amp-hour unit is convenient for simple calculations. For example, a 60-amp-hour battery can supply a current of 1.0 A for 60 h (or 60 A for 1.0 h) before going dead. Find the energy stored in the following. (a) a 58 A-h, 13 V car battery Incorrect: Your answer is incorrect. (b) a 1.3 A-h, 1.6 V flashlight battery
Answer #1

a)

1 hour =3600 s

Energy stored in the battery is

E=VIt=58*13754 KWh

in joules

E=754*3600 J=2.7144*106 Joules

b)

E=1.3*1.6 =2.08 KWh

in joules

E=2.08*3600=7.488*103 Joules